Determine the pH of each of the following solutions.

a.

8.09×10−2 M HClO4

Express your answer to three decimal places.

b.

a solution that is 4.4×10−2 M in HClO4 and 5.2×10−2 M in HCl

Express your answer to two decimal places.

c.

a solution that is 1.04% HCl by mass (Assume a density of 1.01 g/mL for the solution.)

Express your answer to three decimal places.

Thank you!!
Answer #1

a)

HClO4 is a strong acid, hence

[H+] = 8.09 x 10-2 M

pH = -log[H+] = -log[8.09 x 10-2] = 1.09

Answer: pH = 1.09

---------------------------------------------------------------------------------------------------------

b)

HClO4 & HCl both are strong acids, so [H+] are additive

[H+] = 4.4 × 10−2 M + 5.2 × 10−2 M = 9.6 x 10-2 M

pH = -log[H+] = -log[9.6 x 10-2] = 1.02

Answer: pH = 1.02

---------------------------------------------------------------------------------------------------------

c)

1 liter = 1010 g

HCl = 1.04 %

(1.04/100) x 1010 = 10.504 g HCl

moles HCl = 10.504 g / 36.4609 g/mol = 0.28808943279 mol

[H+] = 0.28808943279 M

pH = -log[H+] = -log[0.28808943279] = 0.540

Answer: pH = 0.540 -----------------------> use exact this number including zero to fulfill 3 sf

---------------------------------------------------------------------------------------------------------

Hope this helped you!

Thank You So Much! Please Rate this answer as you wish.("Thumbs Up")